OpenStudy (anonymous):

In circle F, Angle CFD=DFE, mBFA=8x, mAFE=5x+11 and BE and AC are diameters

OpenStudy (anonymous):

find m arc DC

ganeshie8 (ganeshie8):

hint : mBFA + mAFE = 180

OpenStudy (anonymous):

I still dont get it...im just not good at this section

ganeshie8 (ganeshie8):

you're given mBFA and mAFE, right ? plug them in the above equation and gind x first

ganeshie8 (ganeshie8):

mBFA + mAFE = 180 8x + 5x+11 = 180 13x + 11 = 180 13x = 169 x = ?

OpenStudy (anonymous):

Would be 13?

ganeshie8 (ganeshie8):

right, now use that to find m arc DC

ganeshie8 (ganeshie8):

we first find mBFA mBFA = 8x = 8*13 = 104

OpenStudy (anonymous):

yeah

ganeshie8 (ganeshie8):

but, mEFC and mBFA are equal cuz both are vertical angles

OpenStudy (anonymous):

okay

ganeshie8 (ganeshie8):

so mEFC = 104 also, since CFD and DFE are equal, m arc CD = mEFC/2 = 104/2 = ?

OpenStudy (anonymous):

52?

ganeshie8 (ganeshie8):

it has to be !

OpenStudy (anonymous):

Thank you!

ganeshie8 (ganeshie8):

yw !